# MOOSE-Chem: Large Language Models for Rediscovering Unseen Chemistry Scientific Hypotheses

### v4

@@ Appendix A Appendix
A.1 Full Proof / Derivation for the Fundamental Assumption
@@
Here $k\in\mathbb{Z}$ represents the number of inspirations needed for a particular $h$. Typically in chemistry, $k\in[1,3]$.
In other words, given existing knowledge in the background, a majority of chemistry research is about searching knowledge that previously not known to be related to the background but in fact can assist the background, then associate the background knowledge and the searched knowledge in a reasonable way to compose a hypothesis. For example, the proposal of backpropagation can be seen as a hypothesis. In this case, the background knowledge is multi-layer logistic regression, and the searched knowledge is the chain rule in calculus.
Here, we call the searched knowledge as “inspiration”. It is vital that the inspiration should not be known to be related to the background before, or at least should not be used to associate with the background in a known way. Otherwise the hypothesis would not be novel.
Our goal is to transform the seemingly impossible-to-solve $P(h|b)$ into an equivalent form, where each step in the equivalent form is practical and executable. Denoting the full (chemistry) literature as $I$, such that $P(I)=1$. Then a straightforward way of decomposing $P(h|b)$ is by the chain rule based on Equation 4:
$P(h|b)$ $=P(h,i_{1},\ldots,i_{k}|b)$ (5)
@@
=$=\begin{cases}P(h|b,i_{1})\cdot P(i_{1}|b,I)&\text{if }k=1\\ P(h|b,i_{1},\ldots,i_{k})\cdot\prod_{j=2}^{k}P(i_{j}|b,i_{1},\ldots,i_{j-1},I)\cdot P(i_{1}|b,I)&\text{if }k>1\end{cases}$ (7)
Here $I$ is the full inspiration space to search for every single $i$ (here we use the existing chemistry literature, containing up to 3000 papers as $I$). The order of $i_{j}$ is exchangeable.
Equation 7 describes the process of $P(h|b)$ in terms of the knowledge-searching perspective. However, $P(h|b,i_{1},\ldots,i_{k})$ and $P(i_{j}|b,i_{1},\ldots,i_{j-1},I)$ might not be enough practicable, and do not precisely reflect how chemistry researchers find a new $i$. One of the main reasons is that researchers tend to think small step by small step. It would be very challenging to think in terms of a big step without breaking it into several small steps.
@@
To mimic how chemistry researchers conduct research and make it more practicable, we break $P(h|b,i_{1},\ldots,i_{k})$ into a series of recursive smaller steps as
$P(h_{k}|b,i_{1},\ldots,i_{k})$ $\approx P(h_{k}|b,f(b,i_{1},\ldots,i_{k-1}),i_{k})$ if $\text{if }k>1$ (8)
$=P(h_{k}|b,h_{k-1},i_{k})$ if $\text{if }k>1$ (9)
Similarly, we can break $P(i_{j+1}|b,i_{1},\ldots,i_{j},I)$ as
$P(i_{k+1}|b,i_{1},\ldots,i_{k},I)$ $\approx P(i_{k+1}|b,f(b,i_{1},\ldots,i_{k}),I)$ if $\text{if }k>1$ (10)
$=P(i_{k+1}|b,h_{k},I)$ if $\text{if }k>1$ (11)
@@
As a result, to achieve the final $h_{k}$, we need to obtain $h_{1},\ldots,h_{k-1}$ first (if $k>1$). In addition, seeing $h$ as a “state”, and $i$ as an “action”, obtaining $h$ and $i$ through $P(h_{k}|b,h_{k-1},i_{k})$ and $P(i_{k+1}|b,h_{k},I)$ correspondingly indicates a Markov property: (1) a new $h$ only depends on $b$, its previous $h$, and the current $i$; and (2) an $i$ only depends on $b$, $I$, and the previous $h$.
Therefore, if $k>1$,
$P(h|b)$ $=P(i_{1},\ldots,i_{k},h_{1},\ldots,h_{k}|b)$ (12)
$=P(i_{1},h_{1}|b)\cdot P(i_{2},h_{2}|b,i_{1},h_{1})\cdot\ldots\cdot P(i_{k},h_{k}|b,i_{1},\ldots,i_{k-1},h_{1},\ldots,h_{k-1})$ (13)
$\approx P(i_{1},h_{1}|b)\cdot P(i_{2},h_{2}|b,h_{1})\cdot\ldots\cdot P(i_{k},h_{k}|b,h_{k-1})$ (14)
@@
$=P(i_{1}|b,I)\cdot P(h_{1}|b,i_{1})\cdot\prod_{j=1}^{k-1}P(i_{j+1}|b,h_{j},I)\cdot P(h_{j+1}|b,i_{j+1},h_{j})$ (16)
$=\prod_{j=1}^{k}P(i_{j}|b,h_{j-1},I)\cdot P(h_{j}|b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset$ (17)
Although starting from $k>1$, Derivation 17 covers the situation when $k=1$ in Equation 7.
Therefore, in sum, we break up the seemingly impossible question $P(h|b)$ into many practical and executable smaller questions as:
$P(h|b)\approx\prod_{j=1}^{k}P(i_{j}|b,h_{j-1},I)\cdot P(h_{j}|b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset\text{ and }k>=1$ (18)
@@ A.6 More Analysis on EU
Table 15 shows the number of hypotheses receiving high Matched Score from only non-EU branch, only EU branches, and only EU-recombination branch. Here only non-EU branch can be seen as the hypotheses obtained directly without mutations. The hypotheses are from the same experiment in Table 10.
@@ A.12 Prompt to GPT-4o for Matched Score
You are helping to evaluate the quality of a proposed research hypothesis in Chemistry by a phd student. The groundtruth hypothesis will also be provided to compare. Here we mainly focus on whether the proposed hypothesis has covered the key points in terms of the methodology in the groundtruth hypothesis. You will also be given a summary of the key points in the methodology of the groundtruth hypothesis for reference. Please note that for the proposed hypothesis to cover one key point, it is not necessary to explicitly mention the name of the key point, but might also can integrate the key point implicitly in the proposed method. The evaluation criteria is called ’Matched score’, which is in a 6-point Likert scale (from 5 to 0). Particularly, 5 points mean that the proposed hypothesis (1) covers all the key points and leverage them similarly as in the methodology of the groundtruth hypothesis, and (2) does not contain any extra key point that has apparent flaws; 4 points mean that the proposed hypothesis (1) covers all the key points (or at least three key points) and leverage them similarly as in the methodology of the groundtruth hypothesis, (2) but also with extra key points that have apparent flaws; 3 points mean that the proposed hypothesis (1) covers at least two key points and leverage them similarly as in the methodology of the groundtruth hypothesis, (2) but does not cover all key points in the groundtruth hypothesis, (3) might or might not contain extra key points; 2 points mean that the proposed hypothesis (1) covers at least one key point in the methodology of the groundtruth hypothesis, and leverage it similarly as in the methodology of groundtruth hypothesis, (2) but does not cover all key points in the groundtruth hypothesis, and (3) might or might not contain extra key points; 1 point means that the proposed hypothesis (1) covers at least one key point in the methodology of the groundtruth hypothesis, (2) but is used differently as in the methodology of groundtruth hypothesis, and (3) might or might not contain extra key points; 0 point means that the proposed hypothesis does not cover any key point in the methodology of the groundtruth hypothesis at all. Please note that the total number of key points in the groundtruth hypothesis might be less than three, so that multiple points can be given. E.g., there’s only one key point in the groundtruth hypothesis, and the proposed hypothesis covers the one key point, it’s possible to give 2 points, 4 points, and 5 points. In this case, we should choose score from 4 points and 5 points, depending on the existence and quality of extra key points. ’Leveraging a key point similarly as in the methodology of the groundtruth hypothesis’ means that in the proposed hypothesis, the same (or very related) concept (key point) is used in a similar way with a similar goal compared to the groundtruth hypothesis (not necessarily for the proposed hypothesis to be exactly the same with the groudtruth hypothesis to be classified as ’similar’). When judging whether an extra key point has apparent flaws, you should use your own knowledge to judge, but rather than to rely on the count number of pieces of extra key point to judge.

### v5

@@ Appendix A Appendix
A.1 Proof of the Rigorous Decomposition of $P(h\mid b)$ Based on the Fundamental Assumption
@@
Here, $k\in\mathbb{Z}$ represents the number of inspirations needed for a particular $h$. Typically in chemistry, $k\in[1,3]$.
Equation 4 expresses the idea that, for the majority of chemistry hypotheses (if not all), each hypothesis can be formulated as a composition of background knowledge and additional knowledge elements, which we refer to as inspirations. This functional form reflects a universal pattern in hypothesis formulation: regardless of where the inspiration originates—be it prior literature, serendipitous observation, or discussions with peers—the essential step is to identify correct inspirations and integrate them with existing background knowledge in a meaningful way. In this formulation, the process of first collecting and selecting the appropriate background knowledge and then identifying and integrating suitable inspirations constitutes both a necessary and sufficient condition for generating a valid hypothesis $h$.
Here’s an example in chemistry:
  • Research Question: How to obtain $D_{2}$ gas more efficiently?
  • Background Knowledge: The best performing methods are electrocatalytic methods.
  • Inspiration Knowledge 1: Ruthenium as catalyst
  • Inspiration Knowledge 2: Nitrogen-doped electrode
  • Inspiration Knowledge 3: $D_{2}O$ as chemical solution
  • Hypothesis: A nitrogen-doped ruthenium (Ru) electrode can effectively catalyze the reductive deuteration of (hetero)arenes in the presence of $D_{2}O$ in an electrocatalytic method, leading to efficient $D_{2}$ gas production.
Here’s an example in AI:
  • Research Question: How can we automatically update the parameters of a multi-layer logistic regression model using data?
  • Background Knowledge: Multi-layer logistic regression
  • Inspiration Knowledge: The chain rule from calculus
  • Hypothesis: Backpropagation
Here’s another example in AI:
  • Research Question: How can we improve reasoning performance in language models?
  • Background Knowledge: Chain-of-Thought prompting
  • Inspiration Knowledge: Majority voting over multiple reasoning paths
  • Hypothesis: Self-consistency decoding
Here, “background knowledge” and “inspiration knowledge” as illustrated in the example above, can be understood as specific, well-defined knowledge pieces. In practice, however, knowledge retrieval is rarely so clean. Instead of isolating a single knowledge unit, we often retrieve a noisy cluster of information that contains the desired piece along with extraneous content. For instance, when retrieving a paper that includes a relevant inspiration, the paper will inevitably also contain unrelated information that may not be useful for the current research question. Conversely, a single clean inspiration $i$ may be embedded across multiple papers in the literature. This redundancy is beneficial—it increases the likelihood of retrieving $i$ even when searching imperfectly.
In other words, given existing knowledge in the background, a majority of chemistry research is about searching knowledge that previously not known to be related to the background but in fact can assist the background, then associate the background knowledge and the searched knowledge in a reasonable way to compose a hypothesis. Crucially, the inspiration should not be previously known to be related to the background—at least not in a way that has already been used to formulate hypotheses. Otherwise, the resulting hypothesis would lack novelty. This requirement positions the inspiration retrieval task as an inherently out-of-distribution (OOD) problem, where the goal is to surface connections that lie outside established knowledge associations.
Our goal is to transform the seemingly impossible-to-solve $P(h\mid b)$ into an equivalent form, where each step in the equivalent form is practical and executable. Denoting the full inspiration knowledge space as $I$, such that $P(I)=1$. Then a straightforward way of decomposing $P(h\mid b)$ is by the chain rule based on Equation 4:
$P(h\mid b)$ $=P(h,i_{1},\ldots,i_{k}\mid b)$ (5)
@@
=$=\begin{cases}P(h\mid b,i_{1})\cdot P(i_{1}\mid b,I)&\text{if }k=1\\ P(h\mid b,i_{1},\ldots,i_{k})\cdot\prod_{j=2}^{k}P(i_{j}\mid b,i_{1},\ldots,i_{j-1},I)\cdot P(i_{1}\mid b,I)&\text{if }k>1\end{cases}$ (7)
Here, $P(h,i_{1},\ldots,i_{k}\mid b)$ can be interpreted as the joint distribution over $h$ and its associated inspirations conditioned on background knowledge $b$. Equation 5 holds because Equation 4 asserts that $\{i_{1},\ldots,i_{k}\}$ constitutes a necessary and sufficient set of additional knowledge required to compose $h$. $I$ denotes the full inspiration space—that is, the set of all possible knowledge pieces that could serve as an inspiration for generating a new hypothesis. This space includes not only all existing chemistry knowledge but also potentially relevant knowledge from other disciplines that could be leveraged in formulating novel chemistry hypotheses. However, computing over the full space $I$ is computationally infeasible. To make the problem tractable, we approximate $I$ with a large but manageable subset $\hat{I}$, consisting of approximately 3,000 top cited chemistry papers from the existing chemistry literature.
Equation 7 describes the process of $P(h\mid b)$ from a knowledge-searching perspective. However, the terms $P(h\mid b,i_{1},\ldots,i_{k})$ and $P(i_{j}\mid b,i_{1},\ldots,i_{j-1},I)$ may not fully capture how chemistry researchers actually discover new inspirations in practice. One key reason is that researchers typically reason in an incremental fashion, composing hypotheses by integrating one or two knowledge components at a time. It is cognitively and practically difficult to evaluate or integrate all candidate inspirations simultaneously. Instead, researchers iteratively assess partial combinations—gradually building toward a complete hypothesis.
@@
To mimic how chemistry researchers conduct research and make it more practicable, we break $P(h\mid b,i_{1},\ldots,i_{k})$ into a series of recursive smaller steps as
$P(h_{k}\mid b,i_{1},\ldots,i_{k})$ $\approx P(h_{k}\mid b,f(b,i_{1},\ldots,i_{k-1}),i_{k})$ if $\text{if }k>1$ (8)
$=P(h_{k}\mid b,h_{k-1},i_{k})$ if $\text{if }k>1$ (9)
Similarly, we can break $P(i_{j+1}\mid b,i_{1},\ldots,i_{j},I)$ as
$P(i_{k+1}\mid b,i_{1},\ldots,i_{k},I)$ $\approx P(i_{k+1}\mid b,f(b,i_{1},\ldots,i_{k}),I)$ if $\text{if }k>1$ (10)
$=P(i_{k+1}\mid b,h_{k},I)$ if $\text{if }k>1$ (11)
@@
As a result, to achieve the final $h_{k}$, we need to obtain $\{h_{1},\ldots,h_{k-1}\}$ first (if $k>1$). In addition, seeing $h$ as a “state”, and $i$ as an “action”, obtaining $h$ and $i$ through $P(h_{k}\mid b,h_{k-1},i_{k})$ and $P(i_{k+1}\mid b,h_{k},I)$ correspondingly indicates a Markov property: (1) a new $h$ only depends on $b$, its previous $h$, and the current $i$; and (2) an $i$ only depends on $b$, $I$, and the previous $h$.
For brevity, we visualize this as:
$b\xrightarrow{i_{1}}h_{1}\xrightarrow{i_{2}}h_{2}\xrightarrow{\cdots}h_{k-1}\xrightarrow{i_{k}}h_{k}=h,$
where each transition is still conditioned on the background knowledge $b$, though $b$ is omitted from the notation to emphasize the Markov structure of the progression.
Building on this visualization, we formalize the formation of a hypothesis $h$ (specifically, $h=h_{k}$) as a constructive process that sequentially integrates a set of inspirations $\{i_{1},\ldots,i_{k}\}$ into intermediate hypothesis states $\{h_{1},\ldots,h_{k}\}$.
Formally, the conditional probability $P(h\mid b)$ can be expressed as a marginal over all valid sequences of inspirations that can generate $h$:
$P(h\mid b)=\sum_{\pi\in\Pi_{k}}P\big{(}i_{\pi(1)},\ldots,i_{\pi(k)},h_{1},\ldots,h_{k}\mid b\big{)},$ (12)
where $\Pi_{k}$ denotes the set of all permutations of $\{1,\ldots,k\}$ applied to the inspirations $\{i_{1},\ldots,i_{k}\}$ such that the resulting composition yields the final hypothesis $h_{k}=h$, under the assumption in Equation 4, which specifies that $h$ is fully determined by $b$ and $\{i_{1},\ldots,i_{k}\}$. In this context, $h_{j}$ represents the intermediate hypothesis state obtained after integrating $i_{\pi(j)}$ at step $j$.
The degree to which the order of inspirations $\{i_{1},\ldots,i_{k}\}$ affects hypothesis formulation can vary across disciplines. In empirical sciences such as chemistry, the contributions of individual inspirations are largely interchangeable, and their order of integration has limited impact on the final hypothesis, resulting in a large $|\Pi_{k}|$. Conversely, in disciplines like mathematics, where hypotheses (e.g., theorems) often require constructing a specific sequence of lemmas and prior results, the ordering of inspirations is more constrained and may follow a near-deterministic path.
Considering this variance depending on the disciplines, to make this problem tractable and easier to understand, we adopt a fixed constructive path assumption, i.e., $|\Pi_{k}|=1$, selecting a representative order of inspirations $\{i_{1},\ldots,i_{k}\}$.
Under this inspiration-fixed-order assumption, where hypothesis $h$ is constructed through a specific sequence of inspirations $\{i_{1},\ldots,i_{k}\}$, we have:
$P(h\mid b)=P(i_{1},\ldots,i_{k},h_{1},\ldots,h_{k}\mid b),$ (13)
with $h_{j}$ denoting the intermediate hypothesis state after incorporating $i_{j}$, and $h_{k}=h$.
Therefore, if with this inspiration-fixed-order assumption for simplicity, if $k>1$,
$P(h\mid b)$ $=P(i_{1},\ldots,i_{k},h_{1},\ldots,h_{k}\mid b)$ (14)
$=P(i_{1},h_{1}\mid b)\cdot P(i_{2},h_{2}\mid b,i_{1},h_{1})\cdot\ldots\cdot P(i_{k},h_{k}\mid b,i_{1},\ldots,i_{k-1},h_{1},\ldots,h_{k-1})$ (15)
$\approx P(i_{1},h_{1}\mid b)\cdot P(i_{2},h_{2}\mid b,h_{1})\cdot\ldots\cdot P(i_{k},h_{k}\mid b,h_{k-1})$ (16)
@@
$=P(i_{1}\mid b,I)\cdot P(h_{1}\mid b,i_{1})\cdot\prod_{j=1}^{k-1}P(i_{j+1}\mid b,h_{j},I)\cdot P(h_{j+1}\mid b,i_{j+1},h_{j})$ (18)
$=\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset$ (19)
Although starting from $k>1$, Derivation 19 covers the situation when $k=1$ in Equation 7.
Therefore, in sum, we break up the seemingly impossible question $P(h\mid b)$ into many practical and executable smaller questions as:
$P(h\mid b)\approx\prod_{j=1}^{k}P(i_{j}\mid b,h_{j-1},I)\cdot P(h_{j}\mid b,i_{j},h_{j-1}),\text{ where }h_{0}=\emptyset\text{ and }k>=1$ (20)
Of course, without the inspiration-fixed-order assumption, a more complete derivation of $P(h\mid b)$ involves marginalizing over all valid permutations in $\Pi_{k}$:
$P(h\mid b)$ $=\sum_{\pi\in\Pi_{k}}P(i_{\pi(1)},\ldots,i_{\pi(k)},h_{1},\ldots,h_{k}\mid b)$ (21)
$\approx\sum_{\pi\in\Pi_{k}}\prod_{j=1}^{k}P(i_{\pi(j)}\mid b,h_{j-1}^{(\pi)},I)\cdot P(h_{j}^{(\pi)}\mid b,i_{\pi(j)},h_{j-1}^{(\pi)}),$ (22)
where $h_{0}^{(\pi)}=\emptyset$, and $k>=1$. Here, $h_{j}^{(\pi)}$ denotes the intermediate hypothesis state at step $j$ in the permutation $\pi$, which results from sequentially incorporating inspirations in the order $\{i_{\pi(1)},\ldots,i_{\pi(j)}\}$.
@@ A.6 More Analysis on EU
Table 15 shows the number of hypotheses receiving high Matched Score from only non-EU branch, only EU branches, and only EU-recombination branch. Here, only non-EU branch can be seen as the hypotheses obtained directly without mutations. The hypotheses are from the same experiment in Table 10.
@@ A.12 Prompt to GPT-4o for Matched Score
You are helping to evaluate the quality of a proposed research hypothesis in Chemistry by a phd student. The ground truth hypothesis will also be provided to compare. Here, we mainly focus on whether the proposed hypothesis has covered the key points in terms of the methodology in the ground truth hypothesis. You will also be given a summary of the key points in the methodology of the ground truth hypothesis for reference. Please note that for the proposed hypothesis to cover one key point, it is not necessary to explicitly mention the name of the key point, but might also can integrate the key point implicitly in the proposed method. The evaluation criteria is called ’Matched score’, which is in a 6-point Likert scale (from 5 to 0). Particularly, 5 points mean that the proposed hypothesis (1) covers all the key points and leverage them similarly as in the methodology of the ground truth hypothesis, and (2) does not contain any extra key point that has apparent flaws; 4 points mean that the proposed hypothesis (1) covers all the key points (or at least three key points) and leverage them similarly as in the methodology of the ground truth hypothesis, (2) but also with extra key points that have apparent flaws; 3 points mean that the proposed hypothesis (1) covers at least two key points and leverage them similarly as in the methodology of the ground truth hypothesis, (2) but does not cover all key points in the ground truth hypothesis, (3) might or might not contain extra key points; 2 points mean that the proposed hypothesis (1) covers at least one key point in the methodology of the ground truth hypothesis, and leverage it similarly as in the methodology of ground truth hypothesis, (2) but does not cover all key points in the ground truth hypothesis, and (3) might or might not contain extra key points; 1 point means that the proposed hypothesis (1) covers at least one key point in the methodology of the ground truth hypothesis, (2) but is used differently as in the methodology of ground truth hypothesis, and (3) might or might not contain extra key points; 0 point means that the proposed hypothesis does not cover any key point in the methodology of the ground truth hypothesis at all. Please note that the total number of key points in the ground truth hypothesis might be less than three, so that multiple points can be given. E.g., there’s only one key point in the ground truth hypothesis, and the proposed hypothesis covers the one key point, it’s possible to give 2 points, 4 points, and 5 points. In this case, we should choose score from 4 points and 5 points, depending on the existence and quality of extra key points. ’Leveraging a key point similarly as in the methodology of the ground truth hypothesis’ means that in the proposed hypothesis, the same (or very related) concept (key point) is used in a similar way with a similar goal compared to the ground truth hypothesis (not necessarily for the proposed hypothesis to be exactly the same with the groudtruth hypothesis to be classified as ’similar’). When judging whether an extra key point has apparent flaws, you should use your own knowledge to judge, but rather than to rely on the count number of pieces of extra key point to judge.